FIRST MICROWAVE SPECTRA IN DOPED HELIUM CLUSTERS

\underline{I. REINHARD}, C. CALLEGARI, K.K. LEHMANN, and G. SCOLES, Department of Chemistry, Princeton University, Princeton, NJ 08544.

In the framework of the investigation of the relaxation mechanisms of molecular rotational and vibrational excitations in a superfluid helium cluster environment, the first microwave absorption spectra have been obtained for He clusters doped with acetonitrile (CH_{3}CN). Two broad (≈ 1 GHz) peaks centered at about 12.5 and 15 GHz have been observed using bolometric detection of the beam attenuation caused by successive absorption of several microwave photons in a 10 cm long waveguide. The change in rotational constants induced by the presence of the medium preclude assignment based on a mere comparison with gas phase spectra. Further measurements are currently in progress to assign the observed features and determine the rotational relaxation times of the system.